OpenStudy (anonymous):

I dont understand how u get the lcm of 15, 18, and 54

OpenStudy (anonymous):

Just look at the multiples of each one and find which is the first multiple they all have in common. Or you could do the prime factorization and multiply the primes with the largest exponent: 15: 2^0 * 3^1 * 5^1 18: 2^1 * 3^2 * 5^0 54: 2^1 * 3^3 * 5^0 So the lcm(15,18,54) = 2^1 * 3^3 * 5^1 = 270

OpenStudy (anonymous):

thank u